### v1

@@ 4. Trend violation in pre-treatment periods and no misclassification (TVN)
While misclassification is simulated as before, a violation of the PT assumption in pre-treatment periods is simulated by adding a normalized time and group dependent trend: $Y_{g,t}^{TV}(D_{g,t}^{*})=Y_{g,t}(D_{g,t}^{*})+\frac{t\cdot g}{G}$. For simplicity, the treatment effect is kept constant over time now. The bootstrap is carried out as described in Section 5, and we use the 95-th percentile of the bootstrap distribution as critical value. To decrease the computing time we make use of the warp bootstrap method (Giacomini et al., 2013) and only draw a single bootstrap subsample within each Monte Carlo replication, resulting in 2,000 bootstrap samples.
@@ 7 Empirical Illustration
To illustrate the proposed estimators for true and observed switchers as well as the testing procedure, we revisit the application of Berkhout et al. (2024). The authors analyze how the introduction of a computer based testing system designed to curb cheating in high stakes national school examinations affected exam scores in Indonesia. As reported in their paper, cheating was a widespread problem in the nation-wide examinations at the end of junior and senior secondary school. Therefore, the Indonesian government gradually implemented computer based testing (CBT) instead of paper based exams, which made cheating highly unlikely and led to a drop in exam scores. Berkhout et al. (2024) point to anticipatory behavior by schools aware of their future treatment status, as treatment follows voluntary sign-up for CBT. That is, in order to soften the drop in exam scores, schools might try to counteract cheating practices prior to their actual treatment. The empirical findings of Berkhout et al. (2024) support the presence of such anticipatory behavior one year before the CBT implementation.
For the empirical analysis, we follow Berkhout et al. (2024) and use a balanced panel data set on exam results between 2010 and 2019. CBT was introduced between 2015-2019, but the 2015 and 2016 cohorts have been excluded, since their characteristics are different from the comparison groups. The number of observations is thus 38,621 schools for 10 years, see the original paper for further details on the data.
@@
Original $\widehat{\mathrm{DID}}^{S}_{t}$ $\widehat{\mathrm{DID}}^{S^{\ast}}_{t}$ $\widehat{\lambda}_{t-1}$
Cohort 2017 $-5.21^{\ast}$ $-4.63^{\ast}$ / /
Cohort 2018 $-7.45^{\ast}$ $-7.50^{\ast}$ $-7.83^{\ast}$ 0.1580.1580.158
Cohort 2019 $-5.69^{\ast}$ $-5.69^{\ast}$ $-5.91^{\ast}$ 0.2970.2970.297
Combined $-6.30^{\ast}$ $-6.16^{\ast}$ $-7.00^{\ast}$ /
@@
Table 4 summarizes the treatment effect estimates from the original paper of Berkhout et al. (2024, Figure 5 and Table A3) and from our two proposed estimators as well as the estimated anticipation probabilities. Berkhout et al. (2024) use the staggered adoption estimator of Callaway and Sant’Anna (2021) where never treated schools serve as the control group and one year of anticipation is allowed for. We report both the estimates of the cohort specific effects and of the average effects weighted by the number of groups in each cohort. All estimated ATEs are statistically significant at the 5% level. [^13] We construct critical values for $H_{0}:\;ATE=0$ against its negation using the bootstrap as described in Section 5 with $B=999$. The estimates from the ATE estimator for observed switchers are very similar to the ATE estimates in the original study. Conceptually this is to be expected since the estimator of Callaway and Sant’Anna (2021) should also recover the ATE of observed switchers as the bias from anticipation in the treatment group is offset by dropping the first pre-treatment time period, and the bias from anticipation in the comparison group is avoided by using never treated groups as a comparison. In fact, the small numerical differences between our estimator and the estimator of Callaway and Sant’Anna (2021) are likely to stem from the larger comparison group used in our estimator $\widehat{\mathrm{DID}}^{S}_{t}$.
@@
The cohort specific estimates of the ATE of True Switchers from $\widehat{\mathrm{DID}}^{S^{\ast}}_{t}$ on the other hand are somewhat larger in absolute terms than in the original study. [^14] Note that the estimate for $\widehat{\mathrm{DID}}^{S^{\ast}}_{2017}$ is missing, since an already treated comparison group is lacking. Specifically, we see that for the 2018 cohort there is a drop in mean exam scores by 7.83 points, while for the 2019 cohort mean scores are estimated to drop by 5.91 points (scores range from 0 to 100). These somewhat larger decreases relative to the first and second column of Table 4 could capture not only the effect of CBT, but also the effect of anti-cheating measures implemented by schools in anticipation of the start of CBT. Moreover, since the size of the estimated treatment effects decline in magnitude with the post-treatment time periods as reported in Berkhout et al. (2024), it is reasonable that $\widehat{\mathrm{DID}}^{S^{\ast}}_{t}$ are larger in magnitude as they reflect the (average) treatment effect at the time when schools actually switched behavior. Note that the comparison between the combined estimates of the three estimators should be interpreted with caution, since the first two are weighted averages of three cohorts, while the third (estimator for ATE of observed switchers) is a weighted average of two cohorts which display larger treatment effects than the first cohort. Finally, we estimate the anticipation probability to be 15.8% for the 2018 cohort, and 29.7% for the 2019 cohort.
@@
Next and for illustration purposes, the testing framework from Section 5 is applied to test for the presence of parallel pre-trends and anticipatory behavior. We start by carrying out the the tests using never treated schools are used as control groups similar to Berkhout et al. (2024). Moreover, tests are conducted separately for the three treatment cohorts 2017, 2018 and 2019, where we use 499 bootstrap draws to obtain (5%) critical values. Table 5 displays the test results for the null hypotheses of parallel pre-trends, $H_{0}^{PT}$, and of no misclassification (no anticipation), $H_{0}^{MC}$, for each cohort using one comparison time period. [^15] For instance, the PT test for the 2017 cohort compares the outcomes of schools that are treated in 2017 to schools that are never treated by differencing between their 2015 and 2014 outcomes. The 2016 outcome is discarded since it could be affected by anticipatory behavior. The misclassification test then compares the same schools and their outcomes in 2015 and 2016.
@@
In line with Berkhout et al. (2024), $H_{0}^{PT}$ is rejected at the 5% level for the 2017 cohort and not rejected for the 2018 cohort. However, while Berkhout et al. (2024) detect no violation of parallel pre-trends for the 2019 cohort when comparing 2016 and 2017 outcomes, our test rejects the null here $H_{0}^{PT}$ suggesting a possible violation of parallel pre-trends also in that cohort. Similarly, the misclassification tests based on both the sum and the max statistic reject $H_{0}^{MC}$ in the pre-treatment time period for all three cohorts, while the results in Berkhout et al. (2024) only indicated anticipatory effects for the 2018 and 2019 cohorts only. All our test results remain unchanged when using 1% or 10% significance levels instead. Finally, unreported results also show that constructing the test statistics using not yet treated (rather than just never treated) schools as comparison groups lead to the same rejection decisions with one exception: The PT test for the 2018 cohort rejects at the 5% level, but does not reject at the 1% level. This suggests that the PT assumption might be fragile, at least in periods just prior to the anticipation period.
@@ 8 Conclusion
Additionally, we introduce a two-step testing procedure based on moment equality tests to detect the timing and extent of misspecification, providing a diagnostic framework for empirical researchers to assess the validity of DiD designs under staggered adoption.
@@
[…]
@@ 9.1 Appendix: Additional Quantities
To study the asymptotic behavior of the estimator $\widehat{\mathrm{DID}}_{t}^{S^{\ast}}$ in (11), we require some additional notation. To this end, paralleling the notation in De Chaisemartin and D’Haultefoeuille (2020), for $(d,d^{\prime},t)\in\{0,1\}\times\{0,1\}\times\{2,\cdots,T\}$, let
@@
[…]
@@ 9.2 Appendix: Proofs
As remarked in the proof of Theorem 1 in De Chaisemartin and D’Haultefoeuille (2020), it holds by construction that:
@@
[…]
@@
$\quad\ \frac{N_{1,0,t}^{[1]}N_{1,0,t}}{N_{0,0,t}}\mathrm{E}\left[\widehat{\operatorname{DID}}_{t}^{**}\mid\mathbf{D}\right]$
@@
$\quad-\frac{N_{1,0,t}}{N_{0,0,t}}\sum_{g:D_{g,t+1}=1,D_{g,t}=0}\frac{N_{g,t}}{N_{g,t-1}}\sum_{i=1}^{N_{g,t-1}}\mathrm{E}\left[{D}_{i,g,t-1}^{*}\left({Y}_{i,g,t-1}(1)-{Y}_{i,g,t-1}(0)\right)\mid\mathbf{{D}}\right]$
$\quad+\frac{N_{1,0,t}}{N_{0,0,t}}\sum_{g:D_{g,t+1}=1,D_{g,t}=0}N_{g,t}\mathrm{E}\left[{Y}_{1,t}(0)-{Y}_{1,t-1}(0)\right]$
$\quad-\frac{N_{1,0,t}^{[1]}N_{1,0,t}}{N_{0,0,t}^{[1]}N_{0,0,t}}\sum_{g:D_{g,t+1}=0,D_{g,t}=0}\sum_{i=1}^{N_{g,t}}\mathrm{E}\left[{D}_{i,g,t}^{*}\left(Y_{i,g,t}(1)-Y_{i,g,t}(0)\right)\mid\mathbf{D}\right]$
$\quad+\frac{N_{1,0,t}^{[1]}N_{1,0,t}}{N_{0,0,t}^{[1]}N_{0,0,t}}\sum_{g:D_{g,t+1}=0,D_{g,t}=0}\frac{N_{g,t}}{N_{g,t-1}}\sum_{i=1}^{N_{g,t-1}}\mathrm{E}\left[{D}_{i,g,t-1}^{*}\left({Y}_{i,g,t-1}(1)-{Y}_{i,g,t-1}(0)\right)\mid\mathbf{{D}}\right]$
$\quad-\frac{N_{1,0,t}^{[1]}N_{1,0,t}}{N_{0,0,t}^{[1]}N_{0,0,t}}\sum_{g:D_{g,t+1}=0,D_{g,t}=0}N_{g,t}\mathrm{E}\left[{Y}_{1,t}(0)-{Y}_{1,t-1}(0)\right]$
@@
$\quad\ N_{1,0,t}\mathrm{E}\left[\widehat{\operatorname{DID}}_{t}^{*}\mid\mathbf{D}\right]$
@@
$\quad-\sum_{g:D_{g,t}=0,D_{g,t-1}=0}N_{g,t+1}\mathrm{E}\left[Y_{g,t-1}-Y_{g,t-2}\mid\mathbf{D}\right]$
@@
$\quad-\sum_{g:D_{g,t}=1,D_{g,t-1}=0}\sum_{i=1}^{N_{g,t-2}}\mathrm{E}\left[{D}_{i,g,t-2}^{*}\left({Y}_{i,g,t-2}(1)-{Y}_{i,g,t-2}(0)\right)\mid\mathbf{{D}}\right]$
$\quad+\sum_{g:D_{g,t}=1,D_{g,t-1}=0}N_{g,t}\mathrm{E}\left[{Y}_{1,t-1}(0)-{Y}_{1,t-2}(0)\right]$
$\quad-\sum_{g:D_{g,t}=0,D_{g,t-1}=0}\frac{N_{g,t}}{N_{g,t-1}}\sum_{i=1}^{N_{g,t-1}}\mathrm{E}\left[{D}_{i,g,t-1}^{*}\left(Y_{i,g,t-1}(1)-Y_{i,g,t-1}(0)\right)\mid\mathbf{D}\right]$
$\quad+\sum_{g:D_{g,t}=0,D_{g,t-1}=0}\sum_{i=1}^{N_{g,t-2}}\mathrm{E}\left[{D}_{i,g,t-2}^{*}\left({Y}_{i,g,t-2}(1)-{Y}_{i,g,t-2}(0)\right)\mid\mathbf{{D}}\right]$
$\quad-\sum_{g:D_{g,t}=0,D_{g,t-1}=0}N_{g,t}\mathrm{E}\left[{Y}_{1,t-1}(0)-{Y}_{1,t-2}(0)\right]$
@@
$\quad\ \widehat{\mathrm{DID}}^{S^{\ast}}$
@@
$\quad+\frac{\sum_{t\geq 2}\left(1-\widehat{\lambda}_{t-1}\right)\left(P_{1,0,t}\left(\frac{Q_{1,0,t}}{P_{1,0,t}}-\frac{Q_{0,0,t}}{P_{0,0,t}}\right)+P_{1,0,t}\left(\frac{Q_{1,0,t}^{*}}{P_{1,0,t}}-\frac{Q_{0,0,t}^{*}}{P_{0,0,t}}\right)+\frac{P_{1,0,t}^{[1]}P_{1,0,t}}{P_{0,0,t}}\left(\frac{Q_{1,0,t}^{**}}{P_{1,0,t}^{[1]}}-\frac{Q_{0,0,t}^{**}}{P_{0,0,t}^{[1]}}\right)\right)}{\sum_{t\geq 2}\left(\widehat{\lambda}_{t-1}P^{[1]}_{1,0,t-1}+(1-\widehat{\lambda}_{t-1})P_{1,0,t}\right)}$
@@
$\qquad\qquad\left.+\left(1-g_{t}(\boldsymbol{V})\right)\cdot\left(\left(Q_{1,0,t}-\frac{P_{1,0,t}}{P_{0,0,t}}Q_{0,0,t}\right)+\left(Q_{1,0,t}^{*}-\frac{P_{1,0,t}}{P_{0,0,t}}Q_{0,0,t}^{*}\right)+\left(\frac{P_{1,0,t}}{P_{0,0,t}}Q_{1,0,t}^{**}-\frac{P_{1,0,t}^{[1]}P_{1,0,t}}{P_{0,0,t}^{[1]}P_{0,0,t}}Q_{0,0,t}^{**}\right)\right)\right)$
@@
$\quad\ f_{N}\left(\mathrm{E}\left[\mathbf{V}\right]\right)$
@@
$\qquad\qquad+\left(1-g_{t}\left(\mathrm{E}\left[\mathbf{V}\right]\right)\right)\cdot\left(\left(\mathrm{E}\left[Q_{1,0,t}\right]-\frac{\mathrm{E}\left[P_{1,0,t}\right]}{\mathrm{E}\left[P_{0,0,t}\right]}\mathrm{E}\left[Q_{0,0,t}\right]\right)+\left(\mathrm{E}\left[Q_{1,0,t}^{*}\right]-\frac{\mathrm{E}\left[P_{1,0,t}\right]}{\mathrm{E}\left[P_{0,0,t}\right]}\mathrm{E}\left[Q_{0,0,t}^{*}\right]\right)\right.$
$\qquad\qquad\qquad\qquad\qquad\qquad\quad\left.\left.+\left(\frac{\mathrm{E}\left[P_{1,0,t}\right]}{\mathrm{E}\left[P_{0,0,t}\right]}\mathrm{E}\left[Q_{1,0,t}^{**}\right]-\frac{\mathrm{E}\left[P_{1,0,t}^{[1]}\right]\cdot\mathrm{E}\left[P_{1,0,t}\right]}{\mathrm{E}\left[P_{0,0,t}^{[1]}\right]\cdot\mathrm{E}\left[P_{0,0,t}\right]}\mathrm{E}\left[Q_{0,0,t}^{**}\right]\right)\right)\right)$

### v2

@@ 4. Trend violation in pre-treatment periods and no misclassification (TVN)
While misclassification is simulated as before, a violation of the PT assumption in pre-treatment periods is simulated by adding a normalized time and group dependent trend: $Y_{g,t}^{TV}(D_{g,t}^{*})=Y_{g,t}(D_{g,t}^{*})+\frac{t\cdot g}{G}$. For simplicity, the treatment effect is kept constant over time now. The bootstrap is carried out as described in Section 5, and we use the 95-th percentile of the bootstrap distribution as critical value. To decrease the computing time we make use of the warp bootstrap method (Giacomini2013) and only draw a single bootstrap subsample within each Monte Carlo replication, resulting in 2,000 bootstrap samples.
@@ 7 Empirical Illustration
To illustrate the proposed estimators for true and observed switchers as well as the testing procedure, we revisit the application of BPRSS2024. The authors analyze how the introduction of a computer based testing system designed to curb cheating in high stakes national school examinations affected exam scores in Indonesia. As reported in their paper, cheating was a widespread problem in the nation-wide examinations at the end of junior and senior secondary school. Therefore, the Indonesian government gradually implemented computer based testing (CBT) instead of paper based exams, which made cheating highly unlikely and led to a drop in exam scores. BPRSS2024 point to anticipatory behavior by schools aware of their future treatment status, as treatment follows voluntary sign-up for CBT. That is, in order to soften the drop in exam scores, schools might try to counteract cheating practices prior to their actual treatment. The empirical findings of BPRSS2024 support the presence of such anticipatory behavior one year before the CBT implementation.
For the empirical analysis, we follow BPRSS2024 and use a balanced panel data set on exam results between 2010 and 2019. CBT was introduced between 2015-2019, but the 2015 and 2016 cohorts have been excluded, since their characteristics are different from the comparison groups. The number of observations is thus 38,621 schools for 10 years, see the original paper for further details on the data.
@@
Original C&S No Ant. C&H No Ant. $\widehat{\mathrm{DID}}^{S}_{t}$ $\widehat{\mathrm{DID}}^{S^{\ast}}_{t}$ $\widehat{\lambda}_{t-1}$
Cohort 2017 $5.21^{\ast}$ $3.98^{\ast}$ $3.01^{\ast}$ $4.87^{\ast}$ / /
Cohort 2018 $7.45^{\ast}$ $5.61^{\ast}$ $4.48^{\ast}$ $7.16^{\ast}$ $7.83^{\ast}$ $0.158$
Cohort 2019 $5.69^{\ast}$ $^{\ast} 3.78-3.78^{\ast}$ $5.69^{\ast}$ $5.91^{\ast}$ $0.297$
Combined $6.30^{\ast}$ $4.59^{\ast}$ $3.87^{\ast}$ $6.08^{\ast}$ $7.00^{\ast}$ /
@@
Table 4 reports the treatment effect estimates from the original paper of BPRSS2024, two estimators not accounting for anticipation and our two proposed estimators, as well as the estimated anticipation probabilities. BPRSS2024 use the staggered adoption estimator of CSA2021 where never treated schools serve as the control group and one year of anticipation is allowed for (Original). For comparison, we also present the estimator of CSA2021 without accounting for anticipation (C&S No Ant.) and the CH2020 estimator that also does not allow for anticipation (C&H No Ant.). We report both the estimates of the cohort specific effects and of the average effects weighted by the number of groups in each cohort. All estimated ATEs are statistically significant at the 5% level. [^12] We construct critical values for $H_{0}:\;ATE=0$ against its negation using the bootstrap as described in Section 5 with $B=999$.
The estimates from the ATE estimator of observed switchers ($\widehat{\mathrm{DID}}^{S}_{t}$) and from the ATE estimator of True Switchers ($\widehat{\mathrm{DID}}^{S^{\ast}}_{t}$) are substantially larger than those obtained from estimators that do not account for anticipation, underscoring the importance of taking anticipation into account. Specifically, we see for the 2018 cohort a drop in mean exam scores by 7.50 and 7.83 points, respectively (scores range from 0 to 100). In comparison, not accounting for anticipation yields smaller estimated declines of 5.61 (C&S No Ant.) or 4.48 points (C&H No Ant.). The estimates based on CH2020 are even smaller than those from CSA2021 without accounting for anticipation, since the bias not only results from the treated groups, but also from the comparison groups. Additionally, we estimate the anticipation probability ($\widehat{\lambda}_{t-1}$) to be 15.8% for the 2018 cohort, and 29.7% for the 2019 cohort complementing the picture of non-negligible anticipatory behavior among schools.
The estimates from the ATE estimator of observed switchers ($\widehat{\mathrm{DID}}^{S}_{t}$) are very similar to the ATE estimates in the original study. Conceptually, this is to be expected since the estimator of CSA2021 should also recover the ATE of observed switchers as the bias from anticipation in the treatment group is offset by dropping the first pre-treatment time period, and the bias from anticipation in the comparison group is avoided by using never treated groups as a comparison. The small numerical differences between our estimator and the estimator of CSA2021 are likely to stem from the larger comparison group used in our estimator $\widehat{\mathrm{DID}}^{S}_{t}$.
@@
The cohort specific estimates for the ATE of true switchers from $\widehat{\mathrm{DID}}^{S^{\ast}}_{t}$ on the other hand are somewhat larger in absolute terms than in the original study, possibly capturing not only the effect of CBT, but also the effect of anti-cheating measures implemented prior to the start of CBT. [^13] Note that the estimate for $\widehat{\mathrm{DID}}^{S^{\ast}}_{2017}$ is missing, since an already treated comparison group is lacking. Moreover, since the size of the estimated treatment effects decline in magnitude with the post-treatment time periods as reported in BPRSS2024, it is reasonable that $\widehat{\mathrm{DID}}^{S^{\ast}}_{t}$ are larger in magnitude as they reflect the (average) treatment effect at the time when schools actually switched behavior. Note that the combined estimate for the ATE of true switchers should be interpreted with caution, since it is a weighted average of two cohorts only which display larger treatment effects than the first cohort.
@@
Next and for illustration purposes, the testing framework from Section 5 is applied to test for the presence of parallel pre-trends and anticipatory behavior. We start by carrying out the tests using never treated schools as control groups similar to BPRSS2024. Moreover, tests are conducted separately for the three treatment cohorts 2017, 2018 and 2019, where we use 499 bootstrap draws to obtain (5%) critical values. Table 5 displays the test results for the null hypotheses of parallel pre-trends, $H_{0}^{PT}$, and of no misclassification (no anticipation), $H_{0}^{MC}$, for each cohort using one comparison time period. [^14] For instance, the PT test for the 2017 cohort compares the outcomes of schools that are treated in 2017 to schools that are never treated by differencing between their 2015 and 2014 outcomes. The 2016 outcome is discarded since it could be affected by anticipatory behavior. The misclassification test then compares the same schools and their outcomes in 2015 and 2016.
@@
In line with BPRSS2024, $H_{0}^{PT}$ is rejected at the 5% level for the 2017 cohort and not rejected for the 2018 cohort. However, while there is no evidence for a violation of parallel pre-trends in BPRSS2024 for the 2019 cohort, our test rejects the null here $H_{0}^{PT}$ suggesting a possible violation also in that cohort. Similarly, the misclassification tests based on both the sum and the max statistic reject $H_{0}^{MC}$ in the pre-treatment time period for all three cohorts, while the results in BPRSS2024 only indicate anticipatory effects for the 2018 and 2019 cohorts only. All our test results remain unchanged when using 1% or 10% significance levels instead.
Finally, unreported results also show that constructing the test statistics using not yet treated (rather than just never treated) schools as comparison groups lead to the same rejection decisions with one exception: The PT test for the 2018 cohort rejects at the 5% level, but does not reject at the 1% level. This suggests that the PT assumption might be fragile, at least in periods just prior to the anticipation period. Additional evidence for this interpretation comes from PT tests comparing each cohort to never-treated schools and using three comparison periods instead of one [^15] For instance, the PT test for the 2017 cohort compares the outcomes of schools that are treated in 2017 to schools that are never treated by differencing between their 2015 and 2014 outcomes, 2015 and 2013 outcomes, and 2015 and 2012 outcomes.. The results yield rejection decisions across all three cohorts, suggesting increased power when multiple moments are incorporated.
@@ 8 Conclusion
Additionally, we introduce two specification tests based on moment equalities to detect violations of pre-treatment PT and of the timing and extent of misclassification. This provides a diagnostic framework for empirical researchers to assess the validity of DiD designs under staggered adoption.
@@
We illustrate the methods revisiting the introduction of a CBT system designed to curb cheating behavior in high stakes national school examinations on exam scores in Indonesia. The results indicate that a sizable fraction of schools ($\approx 15-30\%$ depending on the cohort) has likely already anticipated measures against cheating in the year prior to the official CBT implementation. Moreover, our estimates for the ATE of observed and actual switching-to-treatment units suggest slightly more pronounced drops in mean exam scores when schools react to cheating for the first time relative to when they have already been exposed to these measures. Our specification tests largely support the evidence for anticipation, though they also hint at a violation of parallel pre-trends in periods just prior to those anticipation periods.
@@ 9.1 Appendix: Additional Quantities
To study the asymptotic behavior of the estimator $\widehat{\mathrm{DID}}_{t}^{S^{\ast}}$ in (11), we require some additional notation. To this end, paralleling the notation in CH2020, for $(d,d^{\prime},t)\in\{0,1\}\times\{0,1\}\times\{2,\cdots,T\}$, let
@@
while $\mathbf{U}$ denotes the vector of stacked $\mathbf{U}(t,l)$ for all combinations $(t,l)$ in $\mathcal{C}_{S}$, which is of dimension $|\mathcal{C}_{S}|\times 1$. We also define the asymptotic variance:
@@ 9.2 Appendix: Proofs
As remarked in the proof of Theorem 1 in CH2020, it holds by construction that:
@@
$\begin{aligned} \mathrm{E}\left[Y_{g,t}-Y_{g,t-1}\mid\mathbf{{D}}\right]&=\frac{1}{N_{g,t}}\sum_{i=1}^{N_{g,t}}\mathrm{E}\left[{D}_{i,g,t}^{*}\left({Y}_{i,g,t}(1)-{Y}_{i,g,t}(0)\right)+{Y}_{i,g,t}(0)\mid\mathbf{{D}}\right]\\ &\quad-\frac{1}{N_{g,t-1}}\sum_{i=1}^{N_{g,t-1}}\mathrm{E}\left[{D}_{i,g,t-1}^{*}\left({Y}_{i,g,t-1}(1)-{Y}_{i,g,t-1}(0)\right)+{Y}_{i,g,t-1}(0)\mid\mathbf{{D}}\right]\\ &=\frac{1}{N_{g,t}}\sum_{i=1}^{N_{g,t}}\mathrm{E}\left[{D}_{i,g,t}^{*}\left({Y}_{i,g,t}(1)-{Y}_{i,g,t}(0)\right)\mid\mathbf{{D}}\right]\\ &\quad-\frac{1}{N_{g,t-1}}\sum_{i=1}^{N_{g,t-1}}\mathrm{E}\left[{D}_{i,g,t-1}^{*}\left({Y}_{i,g,t-1}(1)-{Y}_{i,g,t-1}(0)\right)\mid\mathbf{{D}}\right]\\ &\quad+\mathrm{E}\left[{Y}_{1,t}(0)-{Y}_{1,t-1}(0)\right].\end{aligned}$ (25)
@@
$\qquad\ \frac{N_{1,0,t}^{[1]}N_{1,0,t}}{N_{0,0,t}}\mathrm{E}\left[\widehat{\operatorname{DID}}_{t}^{**}\mid\mathbf{D}\right]$
@@
$\qquad-\frac{N_{1,0,t}}{N_{0,0,t}}\sum_{g:D_{g,t+1}=1,D_{g,t}=0}\frac{N_{g,t}}{N_{g,t-1}}\sum_{i=1}^{N_{g,t-1}}\mathrm{E}\left[{D}_{i,g,t-1}^{*}\left({Y}_{i,g,t-1}(1)-{Y}_{i,g,t-1}(0)\right)\mid\mathbf{{D}}\right]$
$\qquad+\frac{N_{1,0,t}}{N_{0,0,t}}\sum_{g:D_{g,t+1}=1,D_{g,t}=0}N_{g,t}\mathrm{E}\left[{Y}_{1,t}(0)-{Y}_{1,t-1}(0)\right]$
$\qquad-\frac{N_{1,0,t}^{[1]}N_{1,0,t}}{N_{0,0,t}^{[1]}N_{0,0,t}}\sum_{g:D_{g,t+1}=0,D_{g,t}=0}\sum_{i=1}^{N_{g,t}}\mathrm{E}\left[{D}_{i,g,t}^{*}\left(Y_{i,g,t}(1)-Y_{i,g,t}(0)\right)\mid\mathbf{D}\right]$
$\qquad+\frac{N_{1,0,t}^{[1]}N_{1,0,t}}{N_{0,0,t}^{[1]}N_{0,0,t}}\sum_{g:D_{g,t+1}=0,D_{g,t}=0}\frac{N_{g,t}}{N_{g,t-1}}\sum_{i=1}^{N_{g,t-1}}\mathrm{E}\left[{D}_{i,g,t-1}^{*}\left({Y}_{i,g,t-1}(1)-{Y}_{i,g,t-1}(0)\right)\mid\mathbf{{D}}\right]$
$\qquad-\frac{N_{1,0,t}^{[1]}N_{1,0,t}}{N_{0,0,t}^{[1]}N_{0,0,t}}\sum_{g:D_{g,t+1}=0,D_{g,t}=0}N_{g,t}\mathrm{E}\left[{Y}_{1,t}(0)-{Y}_{1,t-1}(0)\right]$
@@
$\qquad\ N_{1,0,t}\mathrm{E}\left[\widehat{\operatorname{DID}}_{t}^{*}\mid\mathbf{D}\right]$
@@
$\qquad-\sum_{g:D_{g,t}=0,D_{g,t-1}=0}N_{g,t+1}\mathrm{E}\left[Y_{g,t-1}-Y_{g,t-2}\mid\mathbf{D}\right]$
@@
$\qquad-\sum_{g:D_{g,t}=1,D_{g,t-1}=0}\sum_{i=1}^{N_{g,t-2}}\mathrm{E}\left[{D}_{i,g,t-2}^{*}\left({Y}_{i,g,t-2}(1)-{Y}_{i,g,t-2}(0)\right)\mid\mathbf{{D}}\right]$
$\qquad+\sum_{g:D_{g,t}=1,D_{g,t-1}=0}N_{g,t}\mathrm{E}\left[{Y}_{1,t-1}(0)-{Y}_{1,t-2}(0)\right]$
$\qquad-\sum_{g:D_{g,t}=0,D_{g,t-1}=0}\frac{N_{g,t}}{N_{g,t-1}}\sum_{i=1}^{N_{g,t-1}}\mathrm{E}\left[{D}_{i,g,t-1}^{*}\left(Y_{i,g,t-1}(1)-Y_{i,g,t-1}(0)\right)\mid\mathbf{D}\right]$
$\qquad+\sum_{g:D_{g,t}=0,D_{g,t-1}=0}\sum_{i=1}^{N_{g,t-2}}\mathrm{E}\left[{D}_{i,g,t-2}^{*}\left({Y}_{i,g,t-2}(1)-{Y}_{i,g,t-2}(0)\right)\mid\mathbf{{D}}\right]$
$\qquad-\sum_{g:D_{g,t}=0,D_{g,t-1}=0}N_{g,t}\mathrm{E}\left[{Y}_{1,t-1}(0)-{Y}_{1,t-2}(0)\right]$
@@
$\qquad\ \widehat{\mathrm{DID}}^{S^{\ast}}$
@@
$\qquad+\frac{\sum_{t\geq 2}\left(1-\widehat{\lambda}_{t-1}\right)\left(P_{1,0,t}\left(\frac{Q_{1,0,t}}{P_{1,0,t}}-\frac{Q_{0,0,t}}{P_{0,0,t}}\right)+P_{1,0,t}\left(\frac{Q_{1,0,t}^{*}}{P_{1,0,t}}-\frac{Q_{0,0,t}^{*}}{P_{0,0,t}}\right)+\frac{P_{1,0,t}^{[1]}P_{1,0,t}}{P_{0,0,t}}\left(\frac{Q_{1,0,t}^{**}}{P_{1,0,t}^{[1]}}-\frac{Q_{0,0,t}^{**}}{P_{0,0,t}^{[1]}}\right)\right)}{\sum_{t\geq 2}\left(\widehat{\lambda}_{t-1}P^{[1]}_{1,0,t-1}+(1-\widehat{\lambda}_{t-1})P_{1,0,t}\right)}$
@@
$\hskip 17.00024pt\hskip 17.00024pt\left.+\left(1-g_{t}(\boldsymbol{V})\right)\cdot\left(\left(Q_{1,0,t}-\frac{P_{1,0,t}}{P_{0,0,t}}Q_{0,0,t}\right)+\left(Q_{1,0,t}^{*}-\frac{P_{1,0,t}}{P_{0,0,t}}Q_{0,0,t}^{*}\right)+\left(\frac{P_{1,0,t}}{P_{0,0,t}}Q_{1,0,t}^{**}-\frac{P_{1,0,t}^{[1]}P_{1,0,t}}{P_{0,0,t}^{[1]}P_{0,0,t}}Q_{0,0,t}^{**}\right)\right)\right)$
@@
$\qquad\ f_{N}\left(\mathrm{E}\left[\mathbf{V}\right]\right)$
@@
$\hskip 18.49988pt\hskip 18.49988pt+\left(1-g_{t}\left(\mathrm{E}\left[\mathbf{V}\right]\right)\right)\cdot\left(\left(\mathrm{E}\left[Q_{1,0,t}\right]-\frac{\mathrm{E}\left[P_{1,0,t}\right]}{\mathrm{E}\left[P_{0,0,t}\right]}\mathrm{E}\left[Q_{0,0,t}\right]\right)+\left(\mathrm{E}\left[Q_{1,0,t}^{*}\right]-\frac{\mathrm{E}\left[P_{1,0,t}\right]}{\mathrm{E}\left[P_{0,0,t}\right]}\mathrm{E}\left[Q_{0,0,t}^{*}\right]\right)\right$.
$\hskip 18.49988pt\hskip 18.49988pt\hskip 18.49988pt\hskip 18.49988pt\hskip 18.49988pt\hskip 18.49988pt\qquad\left.\left.+\left(\frac{\mathrm{E}\left[P_{1,0,t}\right]}{\mathrm{E}\left[P_{0,0,t}\right]}\mathrm{E}\left[Q_{1,0,t}^{**}\right]-\frac{\mathrm{E}\left[P_{1,0,t}^{[1]}\right]\cdot\mathrm{E}\left[P_{1,0,t}\right]}{\mathrm{E}\left[P_{0,0,t}^{[1]}\right]\cdot\mathrm{E}\left[P_{0,0,t}\right]}\mathrm{E}\left[Q_{0,0,t}^{**}\right]\right)\right)\right)$